OpenStudy (anonymous):

A total of 348 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was two times the number of adult tickets sold . How many adult tickets were sold?

OpenStudy (anonymous):

Let Student tickets be S, Let Adult tickets be A. S = 2A total S+A = 348 (2A) +A =348 3A = 348 devide both sides by 3... A = 116 now S = 2 * 116 S = 232